For a $1813 deposit, a bank offers two choices: Account A pays interest at an APR of 6.9%. Account B pays an APR of 3.5%, but pays a one-time bonus of $84 at the end of 2 years. What is the final value of account A minus the final value of account B?
Answer #1

Future value of account A = Present value ( 1+ r)n

Future value of account A = 1,813 ( 1 + 0.069)2

Future value of account A = 1,813 * 1.142761

Future value of account A = $2,071.826

Future value of account B = Present value ( 1+ r)n + 84

Future value of account B = 1,813 ( 1 + 0.035)2 + 84

Future value of account B = 1,813 * 1.071225 + 84

Future value of account B = $2,026.131

A minus the final value of account B = 2,071.826 - 2,026.131

A minus the final value of account B = $45.70